OpenStudy (anonymous):

write this equation in standard form. x squared +3x=-2. show work.

OpenStudy (earthcitizen):

\[x ^{2}+3x+2=0\]

OpenStudy (anonymous):

\[x^2 + 3x = -2 \rightarrow x^2 +3x +2 =0\]